# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 6.2. Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
(6.10) ${\mathbf{U}}^{\text{(cg)},I}_{\infty,\varepsilon}(t;\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\overline{U}^{\text{(cg)}}_{\infty,\varepsilon}(t)\prod\limits_{j\notin I(1)}g_{t}(x^{\prime}_{j}-x_{j}),$
(6.11) ${\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=V^{\text{(cg)}}_{\infty,\varepsilon}(I(1))\prod\limits_{j\notin I(1)}g_{0}(x^{\prime}_{j}-x_{j}),$
where for ${\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime})$, we actually have $\boldsymbol{x}=\boldsymbol{x}^{\prime}$, since we have squeeze the length of V group to 1. To lighten the notations, we combine (6.10) and (6.11) by introducing
(6.12) ${\mathbf{W}}_{\infty,\varepsilon}^{{\rm(cg)},I}(t;\boldsymbol{x},\boldsymbol{x}^{\prime}):=\mathbbm{1}_{\{|I|=3\}}{\mathbf{U}}^{\text{(cg)},I}_{\infty,\varepsilon}(t;\boldsymbol{x},\boldsymbol{x}^{\prime})+\mathbbm{1}_{\{|I|<3\}}{\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime}).$
Finally, we can write down an upper bound for $\limsup_{N\to\infty}M_{N,\varepsilon}^{\varphi,\psi}$ according to summation (1”)-(3”):
(6.13) $\begin{split}&{\mathcal{M}}_{\infty,\varepsilon}^{\varphi,\psi}:=\limsup_{N\to\infty}M_{N,\varepsilon}^{\varphi,\psi}\leq c\|\psi\|_{\infty}^{4}\varepsilon^{3}\sum\limits_{r=1}^{\infty}C^{r}\sum\limits_{I_{1},\cdots,I_{r},|I_{i}|\leq 3,1\leq i\leq r\atop K_{\varepsilon}<s_{1}\leq t_{1}<\cdots<s_{r}\leq t_{r}<s_{r+1}\leq\frac{2}{\varepsilon}}\sum\limits_{\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{r}\in({\mathbbm{Z}})^{4}\atop\boldsymbol{x}^{\prime}_{1},\cdots,\boldsymbol{x}^{\prime}_{r}\in({\mathbbm{Z}})^{4}}{\mathcal{Q}}_{s_{1}}(\varphi_{\varepsilon},\boldsymbol{x}_{1})\\ \times\Big{\{}&\prod\limits_{i=1}^{r-1}{\mathbf{W}}_{\infty,\varepsilon}^{{\rm(cg)},I_{i}}(t_{i}-s_{i};\boldsymbol{x}_{i},\boldsymbol{x}^{\prime}_{i}){\mathcal{Q}}_{(s_{i+1}-t_{i})}(\boldsymbol{x}^{\prime}_{i},\boldsymbol{x}_{i+1})\Big{\}}{\mathbf{W}}_{\infty,\varepsilon}^{{\rm(cg)},I_{r}}(t_{r}-s_{r};\boldsymbol{x}_{r},\boldsymbol{x}^{\prime}_{r}){\mathcal{Q}}_{(s_{r+1}-t_{r})}(\boldsymbol{x}^{\prime}_{r},\mathbbm{1}_{B_{\varepsilon}}).\end{split}$
@@
(6.14) ${\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\sum\limits_{n=1}^{2/\varepsilon}e^{-\lambda n}{\mathcal{Q}}_{n}(\boldsymbol{x},\boldsymbol{x}^{\prime}),$
(6.15) ${\mathbf{U}}_{\lambda,\varepsilon}^{I}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\sum\limits_{n=0}^{2/\varepsilon}e^{-\lambda n}{\mathbf{U}}^{\text{(cg)},I}_{\infty,\varepsilon}(t;\boldsymbol{x},\boldsymbol{x}^{\prime}),$
(6.16) ${\mathbf{V}}_{\lambda,\varepsilon}^{I}(\boldsymbol{x}.\boldsymbol{x}^{\prime})$ $:={\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime}),$
(6.17) ${\mathcal{P}}_{\lambda,\varepsilon}^{I,J}$ :=$:=\begin{cases}{\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}{\mathbf{V}}_{\lambda,\varepsilon}^{J},\quad&\text{if}~{}|J|<3,\\ {\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}{\mathbf{U}}_{\lambda,\varepsilon}^{J},\quad&\text{if}~{}|J|=3.\end{cases}$
@@
(6.18) $\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{x}^{\prime}):=\frac{w_{\varepsilon}^{\otimes 4}(\boldsymbol{x})}{w_{\varepsilon}^{\otimes 4}(\boldsymbol{x}^{\prime})}{\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{x}^{\prime}).$
@@
(6.19) $\Big{|}{\mathcal{M}}_{\infty,\varepsilon}^{\varphi,\psi}\Big{|}\leq ce^{\frac{2\lambda}{\varepsilon}}\|\psi\|_{\infty}^{4}\varepsilon^{3}\sum\limits_{r=1}^{\infty}C^{r}\sum\limits_{I_{1},\cdots,I_{r}\atop|I_{i}|\leq 3,1\leq i\leq r}\Big{\langle}\frac{\varphi_{\varepsilon}^{\otimes 4}}{w_{\varepsilon}^{\otimes 4}},\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{*,I_{1}}\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{I_{1},I_{2}}\cdots\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{I_{r-1},I_{r}}\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I_{r},*}w_{\varepsilon}\mathbbm{1}_{B_{\varepsilon}}^{\otimes 4}\Big{\rangle},$
where “***” is the partition of $\{1,2,3,4\}$ consisting of all singletons, and $\lambda:=\widehat{\lambda}\varepsilon$ with $\widehat{\lambda}$ a large but fixed constant, such that $e^{2\lambda/\varepsilon}$ is bounded. We have the following proposition to control (6.19)
@@
(6.20) $\Big{\|}\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,J}\Big{\|}_{\ell^{q}\to\ell^{q}}$ $\leq c,$
(6.21) $\Big{\|}\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{*,I}\Big{\|}_{\ell^{q}\to\ell^{q}}\leq c\varepsilon^{-\frac{1}{q}},$ $\quad\quad\Big{\|}\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,*}\Big{\|}_{\ell^{q}\to\ell^{q}}\leq c\varepsilon^{-\frac{1}{p}},$
[…]
@@
Recall the definition of $\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{I,J}$ from (6.17). Since we will let $\varepsilon\to 0$, we actually have that
@@
Hence, noting that number of partition of $\{1,2,3,4\}$ is finite, we then have that
@@
(6.24) ${\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{y})\leq\sum\limits_{n=1}^{2/\varepsilon}\prod\limits_{i=1}^{4}g_{n}(y_{i}-x_{i})\leq\begin{cases}\displaystyle\frac{C}{1+|\boldsymbol{x}-\boldsymbol{y}|^{2}},\quad&\text{for all}~{}\boldsymbol{x},\boldsymbol{y}\in({\mathbbm{Z}})^{4},\\[16.0pt] \displaystyle C\varepsilon e^{-\frac{\varepsilon|\boldsymbol{x}-\boldsymbol{y}|^{2}}{C}},\quad&\text{for}~{}|\boldsymbol{x}-\boldsymbol{y}|\geq\frac{1}{\sqrt{\varepsilon}}.\end{cases}$
@@
(6.25) $\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{y}):=\frac{w_{\varepsilon}^{\otimes 4}(\boldsymbol{x})}{w_{\varepsilon}^{\otimes 4}(\boldsymbol{y})}{\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{y})\leq\begin{cases}\displaystyle\frac{Ce^{C_{0}}}{1+|\boldsymbol{x}-\boldsymbol{y}|^{2}}:=\frac{C}{1+|\boldsymbol{x}-\boldsymbol{y}|^{2}},&\quad\text{on}~{}A_{\varepsilon},\\[16.0pt] \displaystyle C\varepsilon e^{-\frac{\varepsilon|\boldsymbol{x}-\boldsymbol{y}|^{2}+C^{\prime}\sqrt{\varepsilon}|\boldsymbol{x}-\boldsymbol{y}|}{C}}:=C_{1}\varepsilon e^{-C_{1}\varepsilon|\boldsymbol{x}-\boldsymbol{y}|},&\quad\text{on}~{}A_{\varepsilon}^{c}.\end{cases}$
To prove (6.20), (6.21) and (6.22), we split the space by $A_{\varepsilon}$ and $A_{\varepsilon}^{c}$, and the computations are exactly the same as those in the proofs for (4.19), (4.20) and (4.21). Therefore, we omit the details. Note that in (4.11), $k$ transition kernels are equipped with ${\mathbbm{E}}[\zeta^{2}]^{k}$, but in (6.5), $r-1$ heat kernels are equipped with $(\sigma_{\varepsilon}^{2})^{r}$. Hence an extra $\sigma_{\varepsilon}^{2}$ is not needed in (6.22), compared to (4.21).
Proof of (6.23).
@@
(6.26) $\sum\limits_{\boldsymbol{x},\boldsymbol{y}\in({\mathbbm{Z}})_{I}^{4}}\varphi(\boldsymbol{x})\widehat{{\mathbf{V}}}_{\infty,\varepsilon}^{I}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq\frac{C}{\left(\log\frac{1}{\varepsilon}\right)^{\frac{3}{4}}}\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}}.$
@@
First note that by Proposition 2.3, $\lim_{N\to\infty}{\mathbbm{E}}[({\mathcal{Z}}_{N})^{2}]$ exists. Hence, $({\mathcal{Z}}_{N})_{N\geq 1}$ are tight and have sequential weak limits. To show the limit is unique, We show that $({\mathcal{Z}}_{N})_{N\geq 1}$ is a Cauchy sequence.
It suffices to show that for any bounded $f:{\mathbbm{R}}\to{\mathbbm{R}}$ with uniformly bounded first three derivatives, the limit $\lim_{N\to\infty}{\mathbbm{E}}[f({\mathcal{Z}}_{N})]$ exists (cf. [Bil13]). We then show that ${\mathbbm{E}}[f({\mathcal{Z}}_{N})]$ is an Cauchy sequence.
@@
Hence, in the following, we work on $\varphi,\psi\in C_{c}({\mathbbm{R}})$.
Note that we have shown that ${\mathcal{Z}}_{N}$ can be approximated by coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon})$ in $L^{2}$. The function $f$ is Lipschitz continuous, and we have that
$\begin{split}\big{|}{\mathbbm{E}}[f({\mathcal{Z}}_{N})]-{\mathbbm{E}}[f(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon}))]\big{|}\leq C\|f^{\prime}\|_{\infty}\big{\|}{\mathcal{Z}}_{N}-\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon})\big{\|}_{L^{2}}=o_{\varepsilon}(1),\quad\text{as}~{}\varepsilon\to 0.\end{split}$
@@
(7.1) $\lim\limits_{\varepsilon\to 0}\limsup_{N\to\infty}\sup_{m,n\geq N}\big{|}{\mathbbm{E}}\big{[}f(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{n,\varepsilon}))\big{]}-{\mathbbm{E}}\big{[}f(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{m,\varepsilon}))\big{]}\big{|}=0.$
To lighten the notations, we write $\mathscr{L}_{\varepsilon}(\Theta_{N})=\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon})$. We will prove (7.1) by a generalized Lindeberg principle (cf. [MOO10, Röl13, CSZ23]). We first verify the local dependence Assumption A.1 for $\Theta(\vec{\mathrm{i}})$. Recall that for indices
@@
(7.2) $A_{\vec{\mathrm{i}}}=\{\vec{\mathrm{j}}=(\mathrm{j},\mathrm{j}^{\prime})\in{\mathbbm{T}}_{\varepsilon}:\{\mathrm{i},\mathrm{i}^{\prime}\}\cap\{\mathrm{j},\mathrm{j}^{\prime}\}\neq\emptyset\}.$

### v3

@@ 6.1 Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
${\mathbf{U}}^{\text{(cg)},I}_{\infty,\varepsilon}(t;\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\overline{U}^{\text{(cg)}}_{\infty,\varepsilon}(t)\prod\limits_{j\notin I(1)}g_{t}(x^{\prime}_{j}-x_{j}),$ (6.12)
${\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=V^{\text{(cg)}}_{\infty,\varepsilon}(I(1))\prod\limits_{j\notin I(1)}g_{0}(x^{\prime}_{j}-x_{j}),$ (6.13)
where for ${\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime})$, we actually have $\boldsymbol{x}=\boldsymbol{x}^{\prime}$, since we have squeezed the length of V group to 1. To lighten the notations, we combine (6.12) and (6.13) by introducing
${\mathbf{W}}_{\infty,\varepsilon}^{{\rm(cg)},I}(t;\boldsymbol{x},\boldsymbol{x}^{\prime}):=\mathbbm{1}_{\{|I|=3\}}{\mathbf{U}}^{\text{(cg)},I}_{\infty,\varepsilon}(t;\boldsymbol{x},\boldsymbol{x}^{\prime})+\mathbbm{1}_{\{|I|<3\}}{\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime}).$ (6.14)
Finally, we can write down an upper bound for ${\mathcal{M}}_{\infty,\varepsilon}^{\varphi,\psi}:=\limsup_{N\to\infty}M_{N,\varepsilon}^{\varphi,\psi}$ according to summation (1”)-(3”):
$\begin{split}{\mathcal{M}}_{\infty,\varepsilon}^{\varphi,\psi}\leq c&\|\psi\|_{\infty}^{4}\varepsilon^{3}\sum\limits_{r=1}^{\infty}C^{r}\!\!\!\!\!\!\sum\limits_{I_{1},\cdots,I_{r},|I_{i}|\leq 3,1\leq i\leq r\atop K_{\varepsilon}<s_{1}\leq t_{1}<\cdots<s_{r}\leq t_{r}<s_{r+1}\leq\frac{2}{\varepsilon}}\sum\limits_{\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{r}\in({\mathbbm{Z}})^{4}\atop\boldsymbol{x}^{\prime}_{1},\cdots,\boldsymbol{x}^{\prime}_{r}\in({\mathbbm{Z}})^{4}}\!\!{\mathcal{Q}}_{s_{1}}(\varphi_{\varepsilon},\boldsymbol{x}_{1}){\mathcal{Q}}_{(s_{r+1}-t_{r})}(\boldsymbol{x}^{\prime}_{r},\mathbbm{1}_{B_{\varepsilon}})\\ \times&\Big\{\prod\limits_{i=1}^{r-1}{\mathbf{W}}_{\infty,\varepsilon}^{{\rm(cg)},I_{i}}(t_{i}-s_{i};\boldsymbol{x}_{i},\boldsymbol{x}^{\prime}_{i}){\mathcal{Q}}_{(s_{i+1}-t_{i})}(\boldsymbol{x}^{\prime}_{i},\boldsymbol{x}_{i+1})\Big\}{\mathbf{W}}_{\infty,\varepsilon}^{{\rm(cg)},I_{r}}(t_{r}-s_{r};\boldsymbol{x}_{r},\boldsymbol{x}^{\prime}_{r}).\end{split}$ (6.15)
@@
${\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\sum\limits_{n=1}^{2/\varepsilon}e^{-\lambda n}{\mathcal{Q}}_{n}(\boldsymbol{x},\boldsymbol{x}^{\prime}),$ (6.16)
${\mathbf{U}}_{\lambda,\varepsilon}^{I}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\sum\limits_{n=0}^{2/\varepsilon}e^{-\lambda n}{\mathbf{U}}^{\text{(cg)},I}_{\infty,\varepsilon}(t;\boldsymbol{x},\boldsymbol{x}^{\prime}),$ (6.17)
${\mathbf{V}}_{\lambda,\varepsilon}^{I}(\boldsymbol{x}.\boldsymbol{x}^{\prime})$ $:={\mathbf{V}}^{\text{(cg)},I}_{\infty,\varepsilon}(\boldsymbol{x},\boldsymbol{x}^{\prime}),$ (6.18)
${\mathcal{P}}_{\lambda,\varepsilon}^{I,J}$ $:=\begin{cases}{\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}{\mathbf{V}}_{\lambda,\varepsilon}^{J},\quad&\text{if}~|J|<3,\\ {\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}{\mathbf{U}}_{\lambda,\varepsilon}^{J},\quad&\text{if}~|J|=3.\end{cases}$ (6.19)
@@
$\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{x}^{\prime}):=\frac{w_{\varepsilon}^{\otimes 4}(\boldsymbol{x})}{w_{\varepsilon}^{\otimes 4}(\boldsymbol{x}^{\prime})}{\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{x}^{\prime}).$ (6.20)
@@
$\Big|{\mathcal{M}}_{\infty,\varepsilon}^{\varphi,\psi}\Big|\leq ce^{\frac{2\lambda}{\varepsilon}}\|\psi\|_{\infty}^{4}\varepsilon^{3}\sum\limits_{r=1}^{\infty}C^{r}\sum\limits_{I_{1},\cdots,I_{r}\atop|I_{i}|\leq 3,1\leq i\leq r}\Big\langle\frac{\varphi_{\varepsilon}^{\otimes 4}}{w_{\varepsilon}^{\otimes 4}},\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{*,I_{1}}\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{I_{1},I_{2}}\cdots\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{I_{r-1},I_{r}}\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I_{r},*}w_{\varepsilon}\mathbbm{1}_{B_{\varepsilon}}^{\otimes 4}\Big\rangle,$ (6.21)
where “*” is the partition of $\{1,2,3,4\}$ consisting of all singletons, and $\lambda:=\widehat{\lambda}\varepsilon$ with $\widehat{\lambda}$ a large but fixed constant, such that $e^{2\lambda/\varepsilon}$ is bounded. We have the following proposition to control (6.21)
@@
$\Big\|\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,J}\Big\|_{\ell^{q}\to\ell^{q}}$ $\leq c,$ (6.22)
$\Big\|\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{*,I}\Big\|_{\ell^{q}\to\ell^{q}}\leq c\varepsilon^{-\frac{1}{q}},$ $\quad\quad\Big\|\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,*}\Big\|_{\ell^{q}\to\ell^{q}}\leq c\varepsilon^{-\frac{1}{p}},$ (6.23)
$\Big\|\widehat{{\mathbf{U}}}_{\widehat{\lambda}\varepsilon,\varepsilon}^{I}\Big\|_{\ell^{q}\to\ell^{q}}$ $\leq\frac{c}{\log\widehat{\lambda}},\quad\text{for}~|I|=3,$ (6.24)
$\Big\|\widehat{{\mathbf{V}}}_{\lambda,\varepsilon}^{I}\Big\|_{\ell^{q}\to\ell^{q}}$ $\leq\frac{c}{(\log\frac{1}{\varepsilon})^{\frac{3}{4}}},\quad\text{for}~|I|\leq 2.$ (6.25)
@@
Recall the definition of $\widehat{{\mathcal{P}}}_{\lambda,\varepsilon}^{I,J}$ from (6.19). Since we will let $\varepsilon\to 0$, we have that
@@
Hence, noting that the number of partitions of $\{1,2,3,4\}$ is finite, we then have that
@@
${\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{y})\leq\sum\limits_{n=1}^{2/\varepsilon}\prod\limits_{i=1}^{4}g_{n}(y_{i}-x_{i})\leq\begin{cases}\displaystyle\frac{C}{1+|\boldsymbol{x}-\boldsymbol{y}|^{2}},\quad&\text{for all}~\boldsymbol{x},\boldsymbol{y}\in({\mathbbm{Z}})^{4},\\[16.0pt] \displaystyle C\varepsilon e^{-\frac{\varepsilon|\boldsymbol{x}-\boldsymbol{y}|^{2}}{C}},\quad&\text{for}~|\boldsymbol{x}-\boldsymbol{y}|\geq\frac{1}{\sqrt{\varepsilon}}.\end{cases}$ (6.26)
@@
$\widehat{{\mathcal{Q}}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{y}):=\frac{w_{\varepsilon}^{\otimes 4}(\boldsymbol{x})}{w_{\varepsilon}^{\otimes 4}(\boldsymbol{y})}{\mathcal{Q}}_{\lambda,\varepsilon}^{I,J}(\boldsymbol{x},\boldsymbol{y})\leq\begin{cases}\displaystyle\frac{Ce^{C_{0}}}{1+|\boldsymbol{x}-\boldsymbol{y}|^{2}}:=\frac{C}{1+|\boldsymbol{x}-\boldsymbol{y}|^{2}},&\text{on}~A_{\varepsilon},\\[16.0pt] \displaystyle C\varepsilon e^{-\frac{\varepsilon|\boldsymbol{x}-\boldsymbol{y}|^{2}+C^{\prime}\sqrt{\varepsilon}|\boldsymbol{x}-\boldsymbol{y}|}{C}}:=C_{1}\varepsilon e^{-C_{1}\varepsilon|\boldsymbol{x}-\boldsymbol{y}|},~~&\text{on}~A_{\varepsilon}^{c}.\end{cases}$ (6.27)
To prove (6.22), (6.23) and (6.24), we work on $A_{\varepsilon}$ and $A_{\varepsilon}^{c}$ separately, and the computations are the same as those in the proof of Proposition 4.3. Therefore, we omit the details. Note that in (4.11), $k$ transition kernels are equipped with ${\mathbbm{E}}[\zeta^{2}]^{k}$, but in (6.5), r1r-1 heat kernels are equipped with $(\sigma_{\varepsilon}^{2})^{r}$. Hence an extra $\sigma_{\varepsilon}^{2}$ is not needed in (6.24), compared to (4.21).
Proof for (6.25).
@@
$\sum\limits_{\boldsymbol{x},\boldsymbol{y}\in({\mathbbm{Z}})_{I}^{4}}\varphi(\boldsymbol{x})\widehat{{\mathbf{V}}}_{\infty,\varepsilon}^{I}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq\frac{C}{\left(\log\frac{1}{\varepsilon}\right)^{\frac{3}{4}}}\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}}.$ (6.28)
@@
First note that by Proposition 2.3, the limit $\lim_{N\to\infty}{\mathbbm{E}}[({\mathcal{Z}}_{N})^{2}]$ exists. Hence, $({\mathcal{Z}}_{N})_{N\geq 1}$ are tight and have sequential weak limits. To show the limit is unique, it suffices to show that for any bounded $f:{\mathbbm{R}}\to{\mathbbm{R}}$ with uniformly bounded first three derivatives, the limit $\lim_{N\to\infty}{\mathbbm{E}}[f({\mathcal{Z}}_{N})]$ exists (cf. [9]). We then show that ${\mathbbm{E}}[f({\mathcal{Z}}_{N})]$ is an Cauchy sequence.
@@
Then by the Lipschitz continuity of $f$, we have that,
$|{\mathbbm{E}}[f({\mathcal{Z}}_{N}(\varphi,\psi))]-{\mathbbm{E}}[f({\mathcal{Z}}_{N}(\varphi,\psi_{\varepsilon}))]|\leq\|f^{\prime}\|_{\infty}\big\|{\mathcal{Z}}_{N}(\varphi,\psi)-{\mathcal{Z}}_{N}(\varphi,\psi_{\varepsilon})\big\|_{2}=o_{\varepsilon}(1).$
Hence, in the following, we only need to work on $\varphi,\psi\in C_{c}({\mathbbm{R}})$.
Note that we have shown that ${\mathcal{Z}}_{N}$ can be approximated by coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon})$ in $L^{2}$. By the Lipschitz continuity of $f$, we have that as $\varepsilon\to 0$,
$\begin{split}\big|{\mathbbm{E}}[f({\mathcal{Z}}_{N})]-{\mathbbm{E}}[f(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon}))]\big|\leq\|f^{\prime}\|_{\infty}\big\|{\mathcal{Z}}_{N}-\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon})\big\|_{L^{2}}=o_{\varepsilon}(1).\end{split}$
@@
$\lim\limits_{\varepsilon\to 0}\limsup_{N\to\infty}\sup_{m,n\geq N}\big|{\mathbbm{E}}\big[f(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{n,\varepsilon}))\big]-{\mathbbm{E}}\big[f(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{m,\varepsilon}))\big]\big|=0.$ (7.1)
To lighten the notations, we write $\mathscr{L}_{\varepsilon}(\Theta_{N})=\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta_{N,\varepsilon})$. We will prove (7.1) by a generalized Lindeberg principle (cf. [31, 34, 15]). We first verify the local dependence Assumption A.1 for $\Theta(\vec{\mathrm{i}})$. Recall that for indices
@@
$A_{\vec{\mathrm{i}}}=\{\vec{\mathrm{j}}=(\mathrm{j},\mathrm{j}^{\prime})\in{\mathbbm{T}}_{\varepsilon}:\{\mathrm{i},\mathrm{i}^{\prime}\}\cap\{\mathrm{j},\mathrm{j}^{\prime}\}\neq\emptyset\}.$ (7.2)